OpenStudy (anonymous):

A pawnbroker buys a tv set and a computer for $600. He sells the comouter at a markup of 30% and the tv set at a markup of 20%. If he makes a profit of $165 on the sale of the two items, hwat did he pay for the computer?

hero (hero):

t + c = 600 1.30t+ 1.20c = 765 Solve the system.

OpenStudy (anonymous):

ooo soo all I was missing was the ones on the %s thought my answer was off

hero (hero):

t(1 + .30) = t + .30t = 1.30t c(1+.20) = c + .20c = 1.20c

hero (hero):

The markups of course is what allows him to make a profit.

OpenStudy (anonymous):

soo is it 1.3t+1.2c=765?

hero (hero):

For the second equation, yes

OpenStudy (anonymous):

first equation is t+c=600 then the second one is 1.3t+1.2c=765

hero (hero):

Correct

OpenStudy (anonymous):

the profit is 165 not 765

OpenStudy (anonymous):

but I know how to solve one equation and pug it into the other=)

hero (hero):

765 is the total amount he will have however.

hero (hero):

If you have $600, then spend it, would reselling and making only $165 back be a profit?

OpenStudy (anonymous):

Ya but they're wanting to know how much he paid for the computer not the total profit

hero (hero):

In order to do that, you have to find the value of variable c. In order to find variable c, you have to set up the proper equations.

hero (hero):

Marie, I don't understand why you are disputing my setup.

OpenStudy (anonymous):

you set it up right I'm just questioning whether it should be 765 or 165

hero (hero):

It goes back to that same thing I was telling you earlier. If you put 165, it would mean he only made 165 after reselling. Making only 165 after reselling is not a profit because 165 < 600 A profit only exists if x > 600 765 > 600, which is a profit of 165 The first equation represents the amount of each item costs equaling their total cost. The second equation represents the re-sell prices equaling the total amount made including the original amount plus the profit.

hero (hero):

There's no other way to do this, sorry.

OpenStudy (anonymous):

7050/11?

hero (hero):

And what might that be?

OpenStudy (anonymous):

vs. 1050/11

OpenStudy (anonymous):

the first fraction is what I got the first time using your way the second one is my way

hero (hero):

FYI, prices are not given in fractions. Put the answer in decimal form up accurate to two decimal places

hero (hero):

That is not correct: Wolfram provides the correct values: http://www.wolframalpha.com/input/?i=t+%2B+c+%3D+600+and+1.30t%2B+1.20c+%3D+765

OpenStudy (anonymous):

okay lol I trust you lol I won't argue next time

OpenStudy (anonymous):

My book says 450

hero (hero):

Yeah, I originally mixed up the variables at the beginning. But 450 and 150 are the answers. If you wanted to nitpick at me, you should have nitpicked me for mixing up the variables :P

hero (hero):

It was supposed to be: c + t = 600 1.30c+ 1.20t = 765 That's the only mistake I made.

hero (hero):

But everything else is right.

OpenStudy (anonymous):

Hey I don't mean to pick on you I just wanna understand this stuff

hero (hero):

I know. Wanna go to vyew?

OpenStudy (anonymous):

?

hero (hero):

Vyew = Collaborative Online Classroom

OpenStudy (anonymous):

Where has this been hiding all this time?

hero (hero):

lol

hero (hero):

It's not a replacement for open study, but it's a great supplement if you want to learn stuff.

OpenStudy (anonymous):

cool

hero (hero):

Click the link ^

OpenStudy (anonymous):

I'm not getting the answer though lol I re-did the variables

hero (hero):

I can show you how

OpenStudy (anonymous):

in the other site?

hero (hero):

Click the link I sent

hero (hero):

Yes

OpenStudy (anonymous):

k

OpenStudy (anonymous):

grait job